# Conformal Tradeoffs: Guarantees Beyond Coverage

### v1

@@ 3.1 Calibrate–and–Audit
Beyond marginal coverage, distribution-free finite-sample statements require an independent audit split (Bian and Barber, 2023; Gibbs et al., 2025). Coverage is special because its split-conformal guarantee is rank-based. For general region–label rates, reusing $\mathcal{D}_{\mathrm{cal}}$ couples region–label indicators to the random threshold and breaks the exchangeable Bernoulli model (Geisser, 1993; Casella and Berger, 2002) needed for exact predictive envelopes; therefore certified envelopes require an independent $\mathcal{D}_{\mathrm{audit}}$ (Appendix D).
@@ 3.2 Certified predictive envelopes for future windows
Section 2 defines the auditable primitive: the joint process $(R_{\tau(\theta)}(X),Y)$ under fixed thresholds, together with KPI indicators obtained as linear projections of the corresponding region–label table (with coverage as a worked example). Under Calibrate–and–Audit, conditional on $\mathcal{D}_{\mathrm{cal}}$ the deployed rule (hence $R_{\tau(\theta)}$ and $\pi$) is fixed, and the audit points are exchangeable with future deployment points. Therefore, for any fixed indicator $I_{\ell}$,
@@
The Beta–Binomial form is not a prior assumption on $p_{\ell}(\theta)$. It is a closed-form expression for the predictive distribution of $K^{m}_{\ell}(\theta)$ induced by the two-stage design: conditional on $\mathcal{D}_{\mathrm{cal}}$, the audit indicators and future-window indicators are i.i.d. Bernoulli trials with a common (unknown) success probability $p_{\ell}(\theta)$. Observing the audit count $K^{\mathrm{audit}}_{\ell}(\theta)$ therefore determines exact finite-sample predictive tails for $K^{m}_{\ell}(\theta)$ under exchangeability, which can be written in Beta–Binomial form.
@@ 3.3 Rate vectors, attainable operational sets, and Pareto filtering
Orientation and Pareto filtering (no cost model).
To summarize negotiation-relevant regimes without committing to a scalar objective, we apply an oriented Pareto filter. Concretely, the analyst chooses, for each displayed KPI, whether it is desirable to be large or small (e.g., lower decisive error exposure is better; higher purity is better). This is an orientation choice, not a pricing model. Formally, fix an orientation vector $s\in\{-1,+1\}^{L}$ and compare points using $s\odot\mathbf{r}(\theta)$. A setting $\theta$ is nondominated if there is no $\theta^{\prime}$ such that
@@
The nondominated set is the negotiation front: any later decision rule that imposes hard constraints or introduces scalar weights must select an operating point on (or near) this front.
@@ 3.4 SSBC as a calibration navigation coordinate
A trade-off map is most useful in practice when it admits a user-facing means of navigation tied to an agreed-upon semantic constraint. In this work we use coverage as that semantic anchor. Coverage is special among the KPIs considered above: under exchangeability, split conformal admits an exact finite-sample rank/Beta characterization of the calibration-conditional coverage (Marques, 2025). SSBC inverts this law to translate a semantic request $(\alpha^{\star},\delta)$ into a concrete discrete choice on the conformal calibration grid (Appendices B and C).
@@
which determine $(\tau_{0},\tau_{1})$ and hence the calibration setting $\theta$. Thus, while the user-facing specification appears four-dimensional $(\alpha_{0}^{\star},\delta_{0},\alpha_{1}^{\star},\delta_{1})$, SSBC collapses it to a two-dimensional navigation coordinate on the calibration grid.
@@
Because $\theta$ is determined by these discrete indices, the calibration family $\Theta$ admits a low-dimensional parameterization (one-dimensional in the pooled case, two-dimensional in the binary Mondrian case) even before accounting for the conservation constraints in the region–label table. Section  4 then makes the resulting coupling explicit in the probability-normalized binary setting and explains the sharp regime boundaries observed when sweeping these navigation indices.
@@ 4 Consequences of a fixed conformal partition in the binary case
Fix a trained scoring function and a realized calibration draw. Calibration selects thresholds $\tau$ and therefore a fixed region process $R_{\tau}(X)$ (Section 2.3). All deployment-facing quantities in this paper are functionals of the joint region–label structure $(R_{\tau}(X),Y)$ (Section 3). In the binary case, treating the conformal interface as the object of planning has three concrete consequences: (i) feasibility is coupled—varying $\tau$ reallocates probability mass across finitely many region types rather than tuning rates independently; (ii) posteriors are piecewise-constant—any region-measurable policy acts only through within-region label composition; and (iii) distribution-free is not cost-agnostic—once actions are wired to regions, coherence with a cost or utility model imposes compatibility constraints that depend on the region–label table (Appendix F). We make these points explicit under probability-normalized scores.
@@ 4.1 Binary conformal geometry and regime boundaries
Sweeping calibration settings moves $\tau$ and reallocates mass across regions. When score support is low-dimensional, small changes in $\tau$ can trigger support discontinuities, and within regimes operational coordinates are coupled. Under probability-normalized scores, swapping $(\tau_{0},\tau_{1})$ exchanges the roles of regions $11$ and $00$. If the deployment policy treats hedging and abstention identically, then this swap does not change action-level behavior, even though set-inclusion coverage changes.
@@ 4.2 Region–label tables, piecewise-constant posteriors, and coherent action
fully characterize the information available to any region-measurable policy. Write $p_{r}:=p_{r,0}+p_{r,1}$ for region mass. Because the interface takes finitely many values, the region-conditioned label composition is constant within each region; downstream evaluation and decision-making therefore reduce to functions of $(p_{r,0},p_{r,1})$, aggregated across regions using the same table $\{p_{r,y}\}$.
Definition (cost-coherence for region-based action).
@@
We say $\tilde{\pi}$ is cost-coherent for the deployed interface if, for every region with $p_{r}>0$, it minimizes conditional risk within that region:
@@
4.3 Conformal partitions are not cost-agnostic under coherent action
Cost-coherence makes a simple point operational: once a conformal partition is deployed, a downstream convention is rational only relative to the information carried by the region label and the costs of available actions. In particular, even if a region is labeled as a singleton under $\pi_{\mathrm{SI}}$, coherence is determined by the region–label composition $(p_{r,0},p_{r,1})$, not by the set label itself.
Reject-option example (Chow-style abstention).
@@ 4.3 Conformal partitions are not cost-agnostic under coherent action
Thus, for a fixed convention, coherence means its prescribed action satisfies the corresponding inequalities in every region (ties allowed). For example, if the policy commits to class 0 in region $10$ (singleton output $\{0\}$ under $\pi_{\mathrm{SI}}$), then coherence requires $p_{10,1}c_{01}\leq p_{10,0}c_{10}$ and $p_{10,1}c_{01}\leq(p_{10,0}+p_{10,1})c_{\mathrm{rej}}$. This also shows why the hard-wired convention “commit on singleton outputs; defer otherwise” is cost-coherent only if, in each singleton region with $p_{r}>0$, the forced commitment has lower expected cost than both rejection and the opposite commitment: a region can be labeled by a singleton set under $\pi_{\mathrm{SI}}$ yet still favor rejection (or even the opposite commitment) for plausible cost ratios. More generally, coherence of any fixed (interface, convention) pair can be expressed as explicit inequalities in the region–label table $\{p_{r,y}\}$ and cost ratios such as $(c_{\mathrm{rej}}/c_{01},c_{\mathrm{rej}}/c_{10})$; see Appendix F. That appendix also gives an inverse view: for a fixed convention, characterize the set of cost ratios under which it is coherent.
@@ 5 Results: validation and deployment-facing planning
This section reports results aligned with the paper’s deployment-facing claims for a single deployed scoring model: the model is treated as fixed infrastructure and we vary only calibration settings (hence the induced region geometry). We focus on three questions: (i) SSBC coverage semantics under calibration randomness and finite deployment windows, (ii) predictive envelopes for deployment-facing operational rates beyond coverage, and (iii) scenario planning via trade-off maps with cost-coherence screening.
@@ 5.1 Numerical Simulations
This subsection validates two core claims: SSBC preserves the intended finite-window coverage semantics under calibration randomness, and Calibrate–and–Audit provides well-centered predictive envelopes for deployment-facing operational rates. Table 1 summarizes semantic coverage behavior, and Figure 4 compares envelope alignment between two-sample and LOO constructions.
@@ 5.1.1 Coverage: numerical realization of SSBC guarantees
This experiment instantiates the exact finite-sample laws underlying SSBC in the finite-window deployment regime, separating calibration randomness from deployment-window noise. Calibration nonconformity scores are drawn i.i.d. from a continuous, heavy-tailed reference distribution, chosen as the absolute Cauchy distribution $|t_{\nu=1}|$ to remain agnostic to model specifics.
@@
Table 1: Representative calibration-conditional coverage violation rates. Target miscoverage $\alpha^{\star}=0.10$, confidence $\delta=0.10$, and deployment window size $m_{\mathrm{infer}}=100$. We report three representative calibration sizes; the complete grid is in Table 7.
$n_{\mathrm{cal}}$ Method $u$ $\alpha_{\mathrm{grid}}$ Obs BetaTheory BBTheory $\alpha_{\mathrm{cont}}$
@@ 5.1.2 Operational rate envelopes: LOO versus two-sample reference
Figure 4: Operational rate envelopes and underlying score distributions. (A) Singleton rate and (B) singleton error for two class prevalences ($p_{\mathrm{class}}=0.10,0.50$) and two class 1 generating distributions ($\mathrm{Beta}(4,3)$ and $\mathrm{Beta}(9,3)$), with class 0 generated by $\mathrm{Beta}(2,7)$. Red rectangles denote the two-sample Calibrate–and–Audit Beta–Binomial ($B$) predictive envelope, and the dashed vertical line is the corresponding $B$ point estimate. Blue and orange markers with horizontal intervals show leave-one-out (LOO) envelopes computed from a single calibration dataset under two inflation levels ($\mathrm{infl}=1,2$). (C1–C4) Histograms of the score geometry shown as predicted class 1 probability (equivalently, nonconformity to class 0) stratified by true class. These distributions make explicit how class prevalence and calibration geometry shape singleton mass and singleton error, and explain the regime-dependent asymmetry of the finite-sample envelopes.
@@ 5.2 Tox21: predictive envelopes under severe class imbalance
Tox21 comprises twelve binary toxicity endpoints (NR and SR pathways). Each endpoint is treated as an independent binary classification problem. Molecules are represented using RDKit (Landrum, 2025) 2D descriptors augmented with Morgan fingerprints (Rogers and Hahn, 2010) (radius $2$, $128$ bits), and classification is performed using CatBoost (Prokhorenkova et al., 2018) with fixed hyperparameters. For each endpoint, data are split randomly into training (50%), calibration (25%), and test (25%), and results are averaged over $100$ splits. Conformal calibration is class-conditional and we compare: (i) standard split conformal, (ii) DKWM correction, and (iii) SSBC, all at $(\alpha,\delta)=(0.10,0.10)$.
@@ 5.2.1 Coverage semantics under small class-conditional calibration sizes
Table 3: Aggregated coverage and prediction-set statistics on Tox21. Results are averaged over twelve endpoints and 100 random splits. Violation rate is the empirical probability that coverage falls below the target level.
Method Coverage Violation Rate Avg. Set Size Singleton Rate
@@
SSBC (proposed) 0.951 0.068 1.54 0.40
5.2.2 Operational envelopes as predictive summaries (LOO versus a two-sample reference)
@@ 5.3 R3: scenario planning on aqueous solubility
This experiment is designed to separate (i) generalization across chemical space from (ii) exchangeable evaluation for conformal and finite-window forecasting. We therefore use scaffold-isolated training, followed by an exchangeable calibration/test design on the held-out scaffold-disjoint pool (Appendix I, dataset/partitioning/model details). Although the base classifier is trained on a three-class discretization of $\log S$ (Insoluble / Moderate / Soluble), deployment-facing planning is posed as a binary campaign objective by merging Moderate and Soluble into a single Soluble class and calibrating prediction sets in $\{\mathrm{Ins},\mathrm{Sol}\}$ (Appendix I, calibration and operationalization).
Operational planning additionally requires a scenario definition: rather than assuming the full calibration pool matches deployment, we restore exchangeability by design for a specified deployment scenario by restricting calibration to a chemically defined subpopulation (Appendix I, Section I.8). In this study the focal scenario is a lipophilic deployment regime $\{\mathrm{MolLogP}>3.5\}$, and SSBC thresholds (and all envelopes derived from them) are computed using the restricted calibration sample.
@@
Figure 5 embodies the main results of this study. The left panel plots the attainable set in a KPI plane spanned by two soluble-class joint rates: irreversible exclusion $P(Y=\mathrm{Sol},C=\{\mathrm{Ins}\})$ and deferral burden $P(Y=\mathrm{Sol},C=\{\mathrm{Sol},\mathrm{Ins}\})$, with color encoding the decisive-correct soluble mass $P(Y=\mathrm{Sol},C=\{\mathrm{Sol}\})$. We highlight nondominated regimes using an orientation-only Pareto filter over these three soluble-class quantities, yielding a shortlist of negotiation-relevant operating regimes.
Because a deployed conformal predictor is an information interface, we also screen regimes for rationality of a fixed downstream wiring given only the conformal output (Section 4.3). The right panel summarizes cost-coherence of the action policy $\{\mathrm{Sol}\}\!\to\!\mathrm{Sol}$, $\{\mathrm{Ins}\}\!\to\!\mathrm{Ins}$, $\{\mathrm{Sol},\mathrm{Ins}\}\!\to\!\mathrm{rej}$ over relative cost ratios $\lambda=c_{01}/c_{10}$ and $\rho=c_{\mathrm{rej}}/c_{10}$. This analysis indicates that points on the Pareto front are cost-coherent for only a limited range of cost ratios: if the cost of irreversible exclusion is too high relative to the cost of deferral, or the cost of deferral is too high relative to the cost of irreversible exclusion, then the action policy yields decisions that yield higher expected costs than the alternative actions in those regions.
@@
Appendix I provides the supporting methodological record and supplementary empirical summaries used to interpret Figure 5: (i) dataset and label construction, scaffold-isolated partitioning, features, and model protocol; (ii) scenario-conditional (tribe-based) calibration restriction and its interpretation; and (iii) diagnostics on the functional roles of $(\alpha_{0},\delta_{0},\alpha_{1},\delta_{1})$ along the Pareto front, together with a discussion of how $\alpha$- and $\delta$-variation act as coarse versus fine adjustments under finite-sample SSBC geometry (Appendix I, Sections I.9I.10).

### v2

@@ 3.1 Calibrate–and–Audit
Beyond marginal coverage, distribution-free finite-sample statements require an independent audit split. Coverage is special because its split-conformal guarantee is rank-based. For general region–class label rates, reusing $\mathcal{D}_{\mathrm{cal}}$ couples region–class label indicators to the random threshold and breaks the exchangeable Bernoulli model (Geisser, 1993; Casella and Berger, 2002) needed for exact predictive envelopes; therefore certified envelopes require an independent $\mathcal{D}_{\mathrm{audit}}$ (Appendix D).
@@ 3.2 Certified predictive envelopes for future windows
Section 2 defined the auditable primitive: the joint process $(R_{\tau(\theta)}(X),Y)$ under fixed thresholds, together with KPI indicators obtained as linear projections of the corresponding region–class label table, with coverage as a worked example. We now use this object to construct predictive envelopes for future operational behavior. Under Calibrate–and–Audit, conditional on $\mathcal{D}_{\mathrm{cal}}$ the deployed rule (hence $R_{\tau(\theta)}$ and $\pi$) is fixed, and the audit points are exchangeable with future deployment points. Therefore, for any fixed indicator $I_{\ell}$,
@@
The Beta–Binomial form is not a prior assumption on $p_{\ell}(\theta)$, but rather a closed-form expression for the predictive distribution of $K^{m}_{\ell}(\theta)$ induced by the two-stage design: conditional on $\mathcal{D}_{\mathrm{cal}}$, the audit indicators and future-window indicators are i.i.d. Bernoulli trials with a common (unknown) success probability $p_{\ell}(\theta)$. Observing the audit count $K^{\mathrm{audit}}_{\ell}(\theta)$ therefore determines exact finite-sample predictive tails for $K^{m}_{\ell}(\theta)$ under exchangeability, which can be written in Beta–Binomial form.
@@ 3.3 Rate vectors, attainable operational sets, and Pareto filtering
Orientation and Pareto filtering
To summarize negotiation-relevant regimes without committing to a scalar objective, we apply an oriented Pareto filter. Concretely, the analyst chooses, for each KPI, whether it is desirable to be large or small (e.g., lower decisive error exposure is better; higher purity is better). This is an orientation choice, not a pricing model. Formally, fix an orientation vector $s\in\{-1,+1\}^{L}$ and compare points using $s\odot\mathbf{r}(\theta)$. A setting $\theta$ is nondominated if there is no $\theta^{\prime}$ such that
@@
The nondominated set is the negotiation front: any later decision rule that imposes hard constraints or introduces scalar weights must select an operating point on (or near) this front. Moreover, if a finalized cost or utility functional is monotone with respect to the chosen orientation, any optimum must be attained on the nondominated front (Miettinen, 1999).
@@ 3.4 SSBC as a calibration navigation coordinate
A trade-off map is most useful in practice when it admits a user-facing means of navigation tied to an agreed-upon semantic constraint. In this work we use coverage as that semantic anchor. Coverage is special among the KPIs considered above: under exchangeability, split conformal admits an exact finite-sample rank/Beta characterization of the calibration-conditional coverage (Marques, 2025; Angelopoulos and Bates, 2023). SSBC inverts this law to translate a semantic request $(\alpha^{\star},\delta)$ into a concrete discrete choice on the conformal calibration grid (Appendices B and C).
@@
which determine $(\tau_{0},\tau_{1})$ and hence the calibration setting $\theta$. Thus, while the user-facing specification appears four-dimensional $(\alpha_{0}^{\star},\delta_{0},\alpha_{1}^{\star},\delta_{1})$, conformal calibration collapses this into a two-dimensional navigation coordinate on the calibration grid.
@@
Because $\theta$ is determined by these discrete indices, the calibration family $\Theta$ admits a low-dimensional parameterization even before accounting for the conservation constraints in the region–class label table. Section  4 then makes the resulting coupling explicit in the probability-normalized binary setting and explains the sharp regime boundaries observed when sweeping these navigation indices.
@@ 4 Consequences of a fixed conformal partition in the binary case
This section examines the geometry induced by a fixed calibrated conformal interface. Given a trained scoring function and a realized calibration draw, calibration selects thresholds $\tau$ and therefore a fixed region process $R_{\tau}(X)$ (Section 2.3). All deployment-facing quantities in this paper are determined by the joint region–class label structure $(R_{\tau}(X),Y)$ (Section 3). In the binary case, this viewpoint has three practical implications. First, feasibility is coupled: varying $\tau$ reallocates probability mass across finitely many region types rather than tuning rates independently. Second, region-based policies depend on within-region class label composition. Third, once actions are tied to regions, coherence with a cost or utility model imposes compatibility constraints that depend on the region–class label table (Appendix F). By coherence we mean that, for each region, the chosen action is optimal among the admissible actions under the given cost structure, so that the resulting expected cost is as small as possible under the corresponding region-conditional label distribution. We make these points explicit under probability-normalized scores.
@@ 4.1 Binary conformal geometry and regime boundaries
Sweeping calibration settings moves $\tau$ and reallocates mass across regions. When score support is low-dimensional, like in the binary case with probability-normalized scores, small changes in $\tau$ can trigger support discontinuities, and within regimes operational coordinates are coupled. Under probability-normalized scores, swapping $(\tau_{0},\tau_{1})$ exchanges the roles of regions $11$ and $00$. If the deployment policy treats hedging and abstention identically, then this swap does not change action-level behavior, even though coverage changes.
@@ 4.2 Region–label tables, piecewise-constant posteriors, and coherent action
fully characterize the information available to any region-associated policy. Write $p_{r}:=p_{r,0}+p_{r,1}$ for region mass. Because the interface takes finitely many values, the region-conditioned label composition is constant within each region; downstream evaluation and decision-making therefore reduce to functions of $(p_{r,0},p_{r,1})$, aggregated across regions using the same table $\{p_{r,y}\}$.
Cost-coherence for region-based action.
@@
We say $\tilde{\pi}$ is cost-coherent for the deployed interface if, for every region with $p_{r}>0$, it minimizes conditional risk within that region. In words, cost coherence means choosing, in each region, an action that minimizes expected cost under the corresponding region-conditional label distribution:
@@
Chow-style Acceptance/Rejection
@@
Thus, for a fixed convention, coherence means its prescribed action satisfies the corresponding inequalities in every region (ties allowed). For example, if the policy commits to class 0 in region $10$ (singleton output $\{0\}$ under $\pi_{\mathrm{SI}}$), then coherence requires $p_{10,1}c_{01}\leq p_{10,0}c_{10}$ and $p_{10,1}c_{01}\leq(p_{10,0}+p_{10,1})c_{\mathrm{rej}}$. This also shows why the hard-wired convention “commit on singleton outputs; defer otherwise” is cost-coherent only if, in each singleton region with $p_{r}>0$, the forced commitment has lower expected cost than both rejection and the opposite commitment: a region can be labeled by a singleton set under $\pi_{\mathrm{SI}}$ yet still favor rejection (or even the opposite commitment) for plausible cost ratios.
This example illustrates the basic point behind cost-coherence: once a conformal partition is deployed, whether a downstream action convention is rational depends on the information carried by the region label together with the costs of the available actions. In particular, even if a region is labeled as a singleton under $\pi_{\mathrm{SI}}$, coherence is determined by the region–label composition $(p_{r,0},p_{r,1})$, not by the set label alone.
More generally, coherence of any fixed (interface, convention) pair can be expressed as explicit inequalities in the region–class label table $\{p_{r,y}\}$ and cost ratios such as $(c_{\mathrm{rej}}/c_{01},c_{\mathrm{rej}}/c_{10})$; see Appendix F. These inequalities restrict the range of cost ratios under which a given convention can be cost-coherent. The appendix develops this inverse view by characterizing, for a fixed convention, the corresponding coherent region of cost ratios.
@@ 5 Results: validation and deployment-facing planning
This section reports results aligned with the paper’s deployment-facing claims for a single deployed scoring model: the model is treated as fixed infrastructure and we vary only calibration settings. We focus on three questions: (i) SSBC coverage semantics under calibration randomness and finite deployment windows, (ii) predictive envelopes for deployment-facing operational rates beyond coverage, and (iii) scenario planning via trade-off maps with cost-coherence screening.
@@ 5.1 Numerical Simulations
This subsection validates two core claims: SSBC preserves the intended finite-window coverage semantics under calibration randomness, and Calibrate–and–Audit provides predictive envelopes for deployment-facing operational rates. Table 1 summarizes semantic coverage behavior, and Figure 4 compares envelope alignment between two-sample and LOO constructions.
@@ 5.1.1 Coverage: numerical realization of SSBC guarantees
This experiment instantiates the exact finite-sample laws underlying SSBC in the finite-window deployment regime, separating calibration randomness from deployment-window noise. Calibration nonconformity scores are drawn i.i.d. from a continuous, heavy-tailed reference distribution, chosen here as the absolute Cauchy distribution $|t_{\nu=1}|$.
@@
Table 1: Representative calibration-conditional coverage violation rates. Target miscoverage $\alpha^{\star}=0.10$, confidence $\delta=0.10$, and deployment window size $m_{\mathrm{infer}}=100$. We report three representative calibration sizes; the complete grid is in Table 7. Obs, Beta and BetaBinom are the observed, theoretical, and Beta-Binomial theoretical violation probabilities, respectively. SSBC achives near-target violation probabilities for all calibration sizes, consistent with Beta-Binomial theory.
$n_{\mathrm{cal}}$ Method $u$ $\alpha_{\mathrm{grid}}$ Obs Beta BetaBinom $\alpha_{\mathrm{cont}}$
@@ 5.1.2 Operational rate envelopes: LOO versus two-sample reference
Figure 4: Operational rate envelopes and underlying score distributions. (A) Singleton rate and (B) singleton error for two class prevalences ($p_{\mathrm{class}}\in\{0.10,0.50\}$) and two class 1 generating distributions ($\mathrm{Beta}(4,3)$ and $\mathrm{Beta}(9,3)$), with class 0 generated by $\mathrm{Beta}(2,7)$. Red rectangles denote the two-sample Calibrate–and–Audit Beta–Binomial ($B$) predictive envelope, and the dashed vertical line is the corresponding $B$ point estimate. Blue and orange markers with horizontal intervals show leave-one-out (LOO) envelopes computed from a single calibration dataset under two inflation levels ($\mathrm{infl}=1,2$). (C1–C4) Histograms of the score geometry shown as predicted class 1 probability (equivalently, nonconformity to class 0) stratified by true class. These distributions make explicit how class prevalence and calibration geometry shape singleton mass and singleton error, and explain the regime-dependent asymmetry of the finite-sample envelopes.
@@ 5.2 Tox21: predictive envelopes under severe class imbalance
Tox21 comprises twelve binary toxicity endpoints (NR and SR pathways). Each endpoint is treated as an independent binary classification problem. Molecules are represented using RDKit (Landrum, 2025) 2D descriptors augmented with Morgan fingerprints (Rogers and Hahn, 2010) (radius $2$, $128$ bits), and classification is performed using CatBoost (Prokhorenkova et al., 2018) with fixed hyperparameters. For each endpoint, data are split randomly into training (50%), calibration (25%), and audit (25%), and results are averaged over $100$ splits. Conformal calibration is class-conditional and we compare: (i) standard split conformal, (ii) DKWM correction, and (iii) SSBC, all at $(\alpha,\delta)=(0.10,0.10)$.
@@ 5.2.1 Coverage semantics under small class-conditional calibration sizes
Table 3: Aggregated coverage and prediction-set statistics on Tox21. Results are averaged over twelve endpoints and 100 random splits. Violation rate is the empirical probability that coverage falls below the target level. Both $\alpha$ and $\delta$ are set to 0.10.
Method Mean Coverage Violation Rate Avg. Set Size Singleton Rate
@@
SSBC 0.951 0.068 1.54 0.40
5.2.2 Operational envelopes as predictive summaries
@@ 5.3 R3: scenario planning on aqueous solubility
This experiment is designed to separate (i) generalization across chemical space from (ii) exchangeable evaluation for conformal and finite-window forecasting. We therefore use a random train/calibration/test split, with the calibration and test sets serving as the exchangeable evaluation design used throughout the solubility experiments (Appendix I, dataset/partitioning/model details). Although the base classifier is trained on a three-class discretization of $\log S$ (Insoluble / Moderate / Soluble), deployment-facing planning is posed as a binary campaign objective by merging Moderate and Soluble into a single Soluble class and calibrating prediction sets in $\{\mathrm{Ins},\mathrm{Sol}\}$ (Appendix I, calibration and operationalization).
For the operational planning demonstration, we simulate a campaign focused on lipophilic compounds. Concretely, we take the focal deployment scenario to be $\{\mathrm{MolLogP}>3.5\}$ and compute SSBC thresholds, together with all derived predictive envelopes, using the calibration sample restricted to this subpopulation (Appendix I, Section I.6).
@@
Figure 5 embodies the main results of this section. The left panel plots the attainable set in a KPI plane spanned by two soluble-class joint rates: irreversible exclusion $P(Y=\mathrm{Sol},C=\{\mathrm{Ins}\})$ and deferral burden $P(Y=\mathrm{Sol},C=\{\mathrm{Sol},\mathrm{Ins}\})$, with color encoding the decisive-correct soluble mass $P(Y=\mathrm{Sol},C=\{\mathrm{Sol}\})$. We highlight nondominated regimes using an orientation-only Pareto filter over these three soluble-class quantities, yielding a shortlist of negotiation-relevant operating regimes.
We also take the inverse pricing view discussed in Section 4.2. For the fixed downstream action convention $\{\mathrm{Sol}\}\!\to\!\mathrm{Sol}$, $\{\mathrm{Ins}\}\!\to\!\mathrm{Ins}$, $\{\mathrm{Sol},\mathrm{Ins}\}\!\to\!\mathrm{rej}$, the right panel characterizes the relative cost ratios $\lambda=c_{01}/c_{10}$ and $\rho=c_{\mathrm{rej}}/c_{10}$ for which this convention is cost-coherent. The result shows that points on the Pareto front are coherent only over a limited region of cost-ratio space: if the cost of irreversible exclusion is too high relative to the cost of deferral, or if the cost of deferral is too high relative to the cost of irreversible exclusion, then the fixed action convention is no longer cost-minimizing in those regions.
@@
Appendix I provides the supporting methodological record and supplementary empirical summaries used to interpret Figure 5: (i) dataset and label construction, random partitioning, features, and model protocol; (ii) scenario-conditional (tribe-based) calibration restriction and its interpretation; and (iii) diagnostics on the functional roles of $(\alpha_{0},\delta_{0},\alpha_{1},\delta_{1})$ along the Pareto front, together with a discussion of how $\alpha$- and $\delta$-variation act as coarse versus fine adjustments under finite-sample SSBC geometry (Appendix I, Sections I.9I.10).